How do you find the x intercept and they y intercept of the graph of the equation? 2y-12x=-6
Found 2 solutions by dlam5, RAY100:
Answer by dlam5(15)

To find "x-intercept" you just cover the y's. So basically, you are doing "opposite.'' Likewise, if you are looking for "y-intercept", you cover up (with your finger or whatever) the x's. Then, you'll find your answer! For any of these LINEAR equations. (note, i caped LINEAR because this process only works for THESE TYPES OF EQUATIONS; (no squares or cubed terms).
Let's solve:
2y-12x=-6
X-INTERCEPT:
cover y's: -12x=-6
x=1/2 (because, -6/-12=1/2)


Y-INTERCEPT:
cover x's: 2y=-6
y=-3 (because, -6/2 = -3)
and there you have it.

:) Thanks

Answer by RAY100(1637)

2y-12x=-6
.
x intercept @ y=o,,,or x= + 1/2,,,,(-12x=-6)
.
y intercept @ x-0,,,or y= -3,,,,,,,,,(2y=-6)
.